how many 2-digit or 3-digits numbers can be formed using the digits 1,2,3,4,5,6,8 and 9 if no repetition is allowed ?



1
Expert's answer
2022-02-09T16:14:20-0500

By Combinatorial Product Rule, the number of 2-digit numbers is equal to "8\cdot 7= 56," the number of 3-dogit numbers is equal to "8\cdot 7\cdot 6=336."

Then the Combinatorial Sum Rule implies that the number of 2-digit or 3-digit numbers is equal to "56+336=392."